OpenStudy (anonymous):

what are the x and y intercepts of f(x)=x^4-6x^2?

OpenStudy (anonymous):

factor first as \[x^2(x^2-6)\] and solve \[x^2(x^2-6)=0\] \[x^2=0\implies x=0\] \[x^2-6=0\implies x^2=6\implies x=\pm\sqrt{6}\]so those are your x intercepts

OpenStudy (anonymous):

y intercept is 0 since \[f(0)=0\]